OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x + 6y = 4 5x + 15y = 10

OpenStudy (anonymous):

A. infinitely many B. 0 C. 2 D. 1

OpenStudy (amoodarya):

divide 1st by 2 divide second by 5

OpenStudy (anonymous):

1, Discriminant = 49 - (4*5*3) = -11 2, nemaproblema: x² - 7/5x = -3/5 (x - 7/10)² - 0.49 = -0.6 (x - 7/10)² = -0.11 Both indicates there is no solution.

OpenStudy (amoodarya):

2x + 6y = 4--->x+3y=2 5x + 15y = 10--->x+3y=2 and now you have 2 variables and one equation second is the same as first

OpenStudy (anonymous):

So the answer is A?

OpenStudy (amoodarya):

can you go on ?

OpenStudy (anonymous):

What do you mean?

OpenStudy (amoodarya):

yes there is infinite solution

OpenStudy (anonymous):

ok Thanks for the help!

OpenStudy (amoodarya):

when variables are more than equations there is infinite solution